# Enhancing Breast Cancer Diagnosis in Mammography: Evaluation and Integration of Convolutional Neural Networks and Explainable AI

### v2

@@ 3 Methodology
This section outlines the comprehensive methodology adopted for the detection and classification of breast cancer using a deep learning model enhanced with Explainable Artificial Intelligence (XAI) techniques. Given the critical nature of breast cancer as a global health issue, the focus is on developing a model that not only achieves high accuracy but also provides transparency in its decision-making process to healthcare professionals.
The methodology is structured into several key components: the dataset used for training and evaluation, the preprocessing steps to prepare the data, the adoption of transfer learning for model development, and the implementation of XAI techniques to explicate the model’s output. Each component is designed to address specific challenges in the task at hand, from handling medical imaging data to enhancing the interpretability of complex deep learning models. By integrating these components, this methodology aims to contribute to the early detection and classification of breast cancer, providing a basis for further research and potential clinical applications as illustrated in Figure 1.
@@ 3.1 Dataset
The Curated Breast Imaging Subset of the Digital Database for Screening Mammography (CBIS-DDSM) is an updated and standardized version of the DDSM, containing 2,620 mammography studies categorized into malignant, benign, and normal classifications. This extensive dataset includes 10,239 images, totalling 163.6 GB, featuring mammograms coupled with accurate pathology annotations through ROI segmentation and bounding boxes. Its comprehensive scale and verified annotations establish CBIS-DDSM as the preferred dataset for this study.
Category Number of Images Comments
Total Images 10,239 -
Malignant Cases 3,000 Includes all modalities
Benign Cases 4,000 Includes all modalities
Normal Cases 3,239 No findings
Image Type Count
Mammograms 10,239 -
Annotations
Mass 5,000
Calcification 2,500
Both 2,000
None 739 Normal cases
@@
Figure 2: A benign mammogram and its Region of Interest (ROI) on the left, contrasted with a malignant mammogram and its ROI on the right [33].
An illustration of this process can be seen in Figure  3, which displays a composite of five ROIs from one mammogram. Accompanying CSV files classify these images into benign and malignant categories, comprising 1,229 benign and 900 malignant cases. Figure  2 exemplifies a benign and a malignant mammogram alongside their respective ROIs from the dataset.
Figure 3: The Five ROI images, proceeding from left to right, merged into a single comprehensive ROI image on the right.
@@ 3.2 Image Pre-processing
Preprocessing is crucial in optimizing the efficacy and computational efficiency of deep learning models applied to medical imaging. For this research, the mammographic images within the dataset were standardized to a resolution of 224x224 pixels. This dimensionality adjustment is followed by a rigorous preprocessing pipeline, meticulously designed to expunge extraneous details while concurrently amplifying diagnostically relevant features within the images. The preprocessing regimen encompasses three principal components: artifact reduction, line removal, and image enhancement, each underpinned by data augmentation strategies to artificially expand the dataset beyond its original volume of 2,129 images.
@@ 3.2.1 Artifact Reduction
The integrity of medical images, specifically mammograms, can be compromised by the presence of non-diagnostic elements such as annotative text and extraneous objects as depicted in Figure 4 and Figure 5. These artifacts, if left unaddressed, could potentially obfuscate the model’s interpretative precision, leading to suboptimal detection and classification performance [33][34]. To mitigate this, artifact removal methods which include morphological opening and contour detection.
@@ 3.2.2 Line Removal
The de-noising through binary masking proved insufficient for the complete removal of all peripheral lines, necessitating further refinement. The intensity thresholding, Gabor filtering, and morphological operations were applied to discern and eliminate residual linear artifacts. This ensures the preservation of breast tissue fidelity while removing non-essential linear distractions  [35][36].
@@ 3.2.3 Image Enhancement
The complexity of mammograms, characterized by the unclear separation between tumorous and dense tissues against a background of fatty tissues, poses a substantial challenge for interpretation based on models. This challenge is intensified in the CBIS-DDSM dataset, where dense tissues are predominantly prevalent. To counteract this, a dual-phase image enhancement strategy was adopted, incorporating Gamma correction followed by the application of Contrast Limited Adaptive Histogram Equalization (CLAHE) - an evolved variant of adaptive histogram equalization [37]. This two-step enhancement, precisely optimized for medical imaging, significantly improves the diagnostic visibility of mammograms, resulting in better model accuracy. The mathematical foundations and detailed processes of CLAHE have been comprehensively explained in [38], providing a solid foundation for its application in enhancing mammographic image quality for deep learning-based diagnostic assessments.
Incorporating these preprocessing techniques within the framework has also been established by prior researchers in [33][34][35][36][37][38], this study not only adheres but also enriches the existing body of knowledge focused on refining mammographic image analysis for enhanced breast cancer detection and classification.
@@ 3.3 Data Augmentation
In the domain of machine learning applications within medical imaging analysis, data augmentation emerges as a critical preprocessing step, especially in tasks involving mammogram classification. The essence of data augmentation lies in its capacity to artificially expand the training dataset through a series of transformations, such as flipping, rotating, and adjusting image properties. This process introduces a broader range of scenarios for the model to learn from, significantly enhancing its robustness and preventing overfitting. It is vital in medical imaging, where the ability to accurately recognize diverse patterns and anomalies can dramatically affect diagnostic outcomes. The goal of implementing specific augmentation strategies is to enrich the dataset with complexities and variabilities reflective of real-world conditions, thereby improving the model’s diagnostic performance.
Before the augmentation process begins, a crucial step involves dividing the collection of mammographic images into training and testing groups, adhering to a 90:10 split. This division is essential for ensuring a balanced representation of both benign and malignant cases in the training and testing phases, maintaining the integrity of the dataset. A balanced representation aids in preventing model bias, ensuring that the diagnostic capabilities of the machine learning model are developed uniformly across different types of cases.
Upon establishing a structured and balanced dataset, the focus shifts to enhancing the robustness and generalizability of the machine learning model. This enhancement is achieved by applying a series of data augmentation techniques to the training set. The deployment of these techniques aims to artificially expand the size of the training dataset by generating modified versions of the images, thereby reducing overfitting and improving the model’s ability to generalize from training to unseen data.
@@
To address this, seven distinct data augmentations were implemented on the training dataset. which include horizontal flipping, vertical flipping, and combined flipping in both directions and rotations (both positive and negative 30 degrees), with and without subsequent flipping as shown in Figure  6. Such comprehensive augmentation ensures the model is exposed to a wide array of image orientations and variations, mirroring the diversity encountered in real-world diagnostic settings.
@@ 3.4 Transfer Learning
The approach involves the use of models pre-trained on large datasets, which are then slightly adjusted or fine-tuned to perform specific tasks, such as identifying abnormalities in mammogram images. Transfer learning is particularly beneficial in situations where annotated medical datasets are scarce. By leveraging knowledge acquired from related tasks, it allows for the efficient training of robust models on smaller datasets, simply by modifying the model’s final layers to adapt to the new task. This technique’s capacity to overcome the challenges posed by limited data availability has been well-documented in the literature of [39]. The ability to transfer learned patterns from one domain to another significantly enhances model performance and accelerates the development process in data-constrained environments. Convolutional Neural Networks (CNNs) are at the forefront of image analysis technologies, with several architectures standing out for their performance. Notable examples include VGG [40], Res-Net [41], Dense-Net [42], Inception [43], and Alex-Net [44]. Each architecture offers distinct advantages, making them suitable for a wide range of applications in deciphering image data. This study aimed to evaluate the performance of various transfer learning models on a dataset curated for mammographic image analysis. The models tested were VGG16, Inception V3, Res-Net18, and Res-Net50. Consistency in dataset usage and data split ratios ensured the reliability of performance comparisons. The objective was to identify the model that The evaluations revealed variations across the tested models, with Res-Net50 emerging as the most effective, achieving a test accuracy of 72%. Enhancement of this model through fine-tuning—a process of re-adjusting pre-trained weights for specific tasks as detailed in [45] led to improved accuracy of 76% as displayed in Table 2.
@@
The fine-tuned Res-Net50 model is instrumental in the integration of XAI within the mammographic analysis.XAI aims to increase transparency in AI decision-making processes, making AI-driven diagnoses more understandable. This is particularly important in the medical sector, where clear diagnostic reasoning can significantly affect patient outcomes. The research findings demonstrate the significant impact of transfer learning on improving the accuracy of AI models for medical image analysis. Additionally, this study lays the foundation for advancements in Explainable AI (XAI), to bridge the gap between AI’s capabilities and human interpretability. This increased understanding has the potential to encourage better collaboration between healthcare professionals and AI systems, ultimately leading to improved patient care.
@@ 4 Results
In the exploration of advanced diagnostic methodologies for breast cancer, the integration of Artificial Intelligence (AI) within mammographic analysis has shown promising potential to augment the precision and efficiency of early detection. This section delves into the empirical findings derived from the application of various interpretability techniques—Grad-CAM, LIME, and SHAP values—on mammographic images. These techniques offer a window into the AI model’s cognitive process, revealing how it discerns between benign and malignant findings. Through a meticulous examination of heatmaps, LIME masks, and SHAP value plots, our study explains the model’s ability to identify and prioritize areas of concern within mammographic images. This not only enhances our understanding of AI-driven diagnostics but also overlays the way for improvements in the interpretability and reliability of AI models in medical imaging.
@@
The application of LIME masks to the original mammograms further refines our understanding of the model’s diagnostic process. By assigning distinct colors to different segments, the LIME analysis dissects the mammogram into regions of varying relevance to the model’s classification decision. This segmentation reveals that the model not only considers the overall tissue density but also pays detailed attention to specific areas that exhibit unusual textural or structural characteristics. For example, Figure 8 showcases how segments marked by LIME correspond to areas where a radiologist might detect signs of pathological changes, thereby validating the model’s accuracy in identifying key diagnostic features.
@@ 5 Evaluation
Explainability methods such as Grad-CAM, LIME, and SHAP are vital for deciphering the workings of deep learning models. However, assessing the safety and reliability of these explanations is equally crucial. Our evaluation, based on the explainability factsheet by Sokol et al. [14], primarily focuses on the fourth dimension: safety. This dimension encompasses four critical factors: information leakage, explanation misuse, explanation invariance and explanation quality. We have systematically evaluated all four of these factors.
@@ 5.1 Information Leakage
To ensure the safety and security of any AI system is important to consider how much knowledge about the model itself is revealed by the information provided in the explanation. An example of information leakage can be the explanations of a k-nearest neighbor model which can reveal the information of the training data points of the model. However, when this factor is considered for our model, it is not a concern as we use deep learning models which are essentially black-box models and require explanations to understand the results. The explanations provided by our chosen methods of Grad-CAM, LIME and SHAP also do not reveal any information about the architecture of the model itself, instead, they help us understand the regions of interest chosen by the model in each sample image. Each output of the models provides visual information which highlights the most important regions in the image.
@@ 5.2 Explanation Misuse
The factor of information misuse comes into consideration based on the information of the model that is leaked via explanations. A model can be stolen or replicated if considerable information about it can be gathered through the explanations, However, this factor does not apply to our model as our methods of explanation do not leak any information regarding the architecture of the model.
@@ 5.3 Explanation Invarience
The third factor of the dimension of safety is explanation invariance. The primary objective of explanations is to develop an understanding of humans for each sample. Hence an explainable system is expected to be stable and consistent.
The factor of stability defines that explanations of the same prediction by a model generated multiple times should result in the same explanation as well. This was evaluated in our model as well by implementing the three XAI techniques multiple times on the same image. An explanation generated by LIME results from the process of random perturbation which generates simulated data in the surroundings of an instance for a simple linear classifier to predict and results in instability as it leads to the generation of different explanations for the same prediction [46]. Hence, while GradCAM and SHAP are stable and produce the same explanation for each prediction, the explanation provided by LIME is unstable, which has been shown in Figure  10.
Figure 10: Multiple LIME results for same mammogram
The factor of consistency defines that explanations of a fixed model for similar data points should also be similar. We have evaluated the consistency of our model by using the ROI data provided in the original CBIS DDSM dataset and checking whether images with similar Regions of Interest have similar XAI results. Figure  11 shows two similar ROI images with similar Grad cam results which shows the consistency of Grad CAM. SHAP works by assigning each feature an importance value for a particular prediction [47]. Hence, similar datapoints are assigned similar importance values which shows the consistency of SHAP. The instability of LIME also makes LIME inconsistent as it generates different explanations for the same prediction, hence comparison of similar data points does not produce similar explanations.
Figure 11: Consistency of Grad CAM results (left column: ROI images, right column: Grad CAM results)
@@ 5.4 Explanation Quality
The final safety measure is the explanation quality which emphasizes the correctness of the explanation provided. While we have achieved the explainability of the model, it is necessary to analyze the correctness of the explanations provided by the different XAI functions to understand the performance of the black-box deep learning model. To analyze the explanations provided by XAI techniques on the performance of our model, the ROI images of the CBIS DDSM dataset have been used to compare the areas of importance that are highlighted by verified pathologists with the areas highlighted by XAI techniques for the deep learning model. The ratio of similarity between the results of XAI and the original ROI images indicated how well the deep learning model is detecting the correct parts of the images for its classification. To create this ratio, we have created binary masks of the explanations provided by GradCAM and LIME which are shown in Figure  12.
@@
Figure  13 shows the original mammogram, the pre-processed image, the Grad CAM output, the mask generated from the Grad CAM heatmap and the original ROI from the dataset.
@@
Figure  14 shows the original mammogram, the pre-processed image, the LIME output, the mask generated from the LIME segments and the original ROI from the dataset.
The binary masks are then compared to the ROI images using Hausdorff distance. Hausdroff distance is a common measure for comparing the dissimilarity between image segmentation and point sets [48]. This method is commonly used in the medical domain to evaluate segmentation in medical images such as computed tomography (CT) and magnetic resonance (MRI) [49]. The Hausdorff distance $\check{H}$ between two point sets $A$ and $B$ is the maximum distance between each point as shown in the following equation [48].
@@
When comparing the ROI images with the masks produced by GradCAM, the Hausdorff distance values of our dataset range from 1 to 133 with the mean Hausdorff value being 18.
When comparing the ROI images with the masks produced by GradCAM, the Hausdorff distance values of our dataset range from 1 to 160 with the mean Hausdorff value being 86. This indicates that the performance of the Grad-CAM explainability technique is better than the explainability provided by LIME.

### v3

@@ 3 Methodology
This section outlines the comprehensive methodology adopted for the detection and classification of breast cancer using a Neural networks with XAI techniques. Given the critical nature of breast cancer as a global health issue, the focus is on developing a model that not only achieves high accuracy but also provides transparency in its decision-making process to healthcare professionals.
The methodology is structured into several key components: the dataset used for training and evaluation, the preprocessing steps to prepare the data, the adoption of transfer learning for model development, and the implementation of XAI techniques to explicate the model’s output. Each component is designed to address specific challenges in the task at hand, from handling medical imaging data to enhancing the interpretability of complex DL models. By integrating these components, this methodology aims to contribute to the early detection and classification of breast cancer, providing a basis for further research and potential clinical applications, as illustrated in Figure 1.
@@ 3.1 Dataset
The Curated Breast Imaging Subset of the Digital Database for Screening Mammography (CBIS-DDSM) is an updated and standardized version of the DDSM, containing 2,620 mammography studies categorized into malignant, benign, and normal classifications as depicted in Table 1 . This extensive dataset includes 10,239 images, totaling 163.6 GB, featuring mammograms coupled with accurate pathology annotations through ROI segmentation and bounding boxes. Its comprehensive scale and verified annotations establish CBIS-DDSM as the preferred dataset for this study.
Characteristic Total Count Categories Origin
Total Images 10,239 Malignant, Benign, Normal USA
Total Subjects 6,671
@@
Figure 2: A benign mammogram and its Region of Interest (ROI) on the left are contrasted with a malignant mammogram and its ROI on the right [33].
An illustration of this process can be seen in Figure 3, which displays a composite of five ROIs from one mammogram. Accompanying CSV files classify these images into benign and malignant categories, comprising 1,229 benign and 900 malignant cases. Figure 2 exemplifies a benign and a malignant mammogram alongside their respective ROIs from the dataset.
Figure 3: The five ROI images, proceeding from left to right, are merged into a single comprehensive ROI image on the right.
@@ 3.2 Image Pre-processing
Preprocessing is crucial in optimizing the efficacy and computational efficiency of DL models applied to medical imaging. For this research, the mammographic images within the dataset were standardized to a resolution of 224x224 pixels. This dimensionality adjustment is followed by a rigorous preprocessing pipeline meticulously designed to expunge extraneous details while simultaneously amplifying diagnostically relevant features within the images. The preprocessing regimen encompasses three principal components: artifact reduction, line removal, and image enhancement, each underpinned by data augmentation strategies to expand the dataset beyond its original volume of 2,129 images.
@@ 3.2.1 Artifact Reduction
The integrity of medical images, specifically mammograms, can be compromised by the presence of non-diagnostic elements such as annotative text and extraneous objects, as depicted in Figures 4 and 5. These artifacts, if left unaddressed, could potentially obfuscate the model’s interpretative precision, leading to suboptimal detection and classification performance [33][34]. To mitigate this, artifact removal methods, which include morphological opening and contour detection,.
@@ 3.2.2 Line Removal
The de-noising through binary masking proved insufficient for the complete removal of all peripheral lines, necessitating further refinement. Intensity thresholding, Gabor filtering, and morphological operations were applied to discern and eliminate residual linear artifacts. This ensures the preservation of breast tissue fidelity while removing non-essential linear distractions.  [35][36].
@@ 3.2.3 Image Enhancement
The complexity of mammograms, characterized by the unclear separation between tumorous and dense tissues against a background of fatty tissues, poses a substantial challenge for interpretation based on models. This challenge is intensified in the CBIS-DDSM dataset, where dense tissues are predominantly prevalent. To counteract this, a dual-phase image enhancement strategy was adopted, incorporating Gamma correction followed by the application of Contrast Limited Adaptive Histogram Equalization (CLAHE), an evolved variant of adaptive histogram equalization [37]. This two-step enhancement, precisely optimized for medical imaging, significantly improves the diagnostic visibility of mammograms, resulting in better model accuracy. The mathematical foundations and detailed processes of CLAHE have been comprehensively explained in [38], providing a solid foundation for its application in enhancing mammographic image quality for DL-based diagnostic assessments.
Incorporating these preprocessing techniques within the framework has also been established by prior researchers in [33][34][35][36][37][38]. This study not only adheres but also enriches the existing body of knowledge focused on refining mammographic image analysis for enhanced breast cancer detection and classification.
@@ 3.3 Data Augmentation
In the domain of ML applications within medical imaging analysis, data augmentation emerges as a critical preprocessing step, especially in tasks involving mammogram classification. The essence of data augmentation lies in its capacity to augment the training dataset through a series of transformations, such as flipping, rotating, and adjusting image properties. This process introduces a broader range of scenarios for the model to learn from, significantly enhancing its robustness and preventing overfitting. It is vital in medical imaging, where the ability to accurately recognize diverse patterns and anomalies can dramatically affect diagnostic outcomes. The goal of implementing specific augmentation strategies is to enrich the dataset with complexities and variabilities reflective of real-world conditions, thereby improving the model’s diagnostic performance.
Before the augmentation process begins, a crucial step involves dividing the collection of mammographic images into training and testing groups, adhering to a 90:10 split. This division is essential for ensuring a balanced representation of both benign and malignant cases in the training and testing phases and maintaining the integrity of the dataset. A balanced representation aids in preventing model bias, ensuring that the diagnostic capabilities of the ML model are developed uniformly across different types of cases.
Upon establishing a structured and balanced dataset, the focus shifts to enhancing the robustness and generalizability of the model. This enhancement is achieved by applying a series of data augmentation techniques to the training set. The deployment of these techniques aims to increase the size of the training dataset by generating modified versions of the images, thereby reducing overfitting and improving the model’s ability to generalize from training to unseen data.
@@
To address this, seven distinct data augmentations were implemented on the training dataset. which include horizontal flipping, vertical flipping, and combined flipping in both directions and rotations (both positive and negative 30 degrees), with and without subsequent flipping as shown in Figure 6. Such comprehensive augmentation ensures the model is exposed to a wide array of image orientations and variations, mirroring the diversity encountered in real-world diagnostic settings.
@@ 3.4 Transfer Learning
The approach involves the use of models pre-trained on large datasets, which are then slightly adjusted or fine-tuned to perform specific tasks, such as identifying abnormalities in mammogram images. Transfer learning is particularly beneficial in situations where annotated medical datasets are scarce. By leveraging knowledge acquired from related tasks, it allows for the efficient training of robust models on smaller datasets, simply by modifying the model’s final layers to adapt to the new task. This technique’s capacity to overcome the challenges posed by limited data availability has been well documented in the literature of [39]. The ability to transfer learned patterns from one domain to another significantly enhances model performance and accelerates the development process in data-constrained environments. Convolutional neural Networks (CNNs) are at the forefront of image analysis technologies, with several architectures standing out for their performance. Notable examples include VGG [40], Res-Net [41], Dense-Net [42], Inception [43], and Alex-Net [44]. Each architecture offers distinct advantages, making them suitable for a wide range of applications in deciphering image data. This study aimed to evaluate the performance of various transfer learning models on a dataset curated for mammographic image analysis. The models tested were VGG16, Inception V3, Res-Net18, and Res-Net50. Consistency in dataset usage and data split ratios ensured the reliability of performance comparisons. The objective was to identify the model that The evaluations revealed variations across the tested models, with Res-Net50 emerging as the most effective, achieving a test accuracy of 72%. Enhancement of this model through fine-tuning—a process of re-adjusting pre-trained weights for specific tasks as detailed in [45]—led to improved accuracy of 76% as displayed in Table 2.
@@
The fine-tuned Res-Net50 model is instrumental in the integration of XAI within the mammographic analysis. XAI aims to increase transparency in AI decision-making processes, making AI-driven diagnoses more understandable. This is particularly important in the medical sector, where clear diagnostic reasoning can significantly affect patient outcomes. The research findings demonstrate the significant impact of transfer learning on improving the accuracy of AI models for medical image analysis. Additionally, this study lays the foundation for advancements in Explainable AI (XAI) to bridge the gap between AI’s capabilities and human interpretability. This increased understanding has the potential to encourage better collaboration between healthcare professionals and AI systems, ultimately leading to improved patient care.
@@ 4 Results
In the exploration of advanced diagnostic methodologies for breast cancer, the integration of AI within mammographic analysis has shown promising potential to augment the precision and efficiency of early detection. This section delves into the empirical findings derived from the application of various interpretability techniques— Grad-CAM, LIME, and SHAP values—on mammographic images. These techniques offer a window into the AI model’s cognitive process, revealing how it discerns between benign and malignant findings. Through a meticulous examination of heatmaps, LIME masks, and SHAP value plots, our study explains the model’s ability to identify and prioritize areas of concern within mammographic images. This not only enhances our understanding of AI-driven diagnostics but also paves the way for improvements in the interpretability and reliability of AI models in medical imaging.
@@
The application of LIME masks to the original mammograms further refines our understanding of the model’s diagnostic process. By assigning distinct colors to different segments, the LIME analysis dissects the mammogram into regions of varying relevance to the model’s classification decision. This segmentation reveals that the model not only considers the overall tissue density but also pays detailed attention to specific areas that exhibit unusual textural or structural characteristics. For example, Figure 8 shows how segments marked by LIME correspond to areas where a radiologist might detect signs of pathological changes, thereby validating the model’s accuracy in identifying key diagnostic features.
@@ 5 Evaluation
Explainability methods such as Grad-CAM, LIME, and SHAP are vital for deciphering the workings of DL models. However, assessing the safety and reliability of these explanations is equally crucial. Our evaluation, based on the explainability factsheet by Sokol et al. [14], primarily focuses on the fourth dimension: safety. This dimension encompasses four critical factors: information leakage, explanation misuse, explanation invariance, and explanation quality. We have systematically evaluated all four of these factors.
@@ 5.1 Information Leakage
To ensure the safety and security of any AI system, it is important to consider how much knowledge about the model itself is revealed by the information provided in the explanation. An example of information leakage can be the explanations of a k-nearest neighbor model, which can reveal the information of the training data points of the model. However, when this factor is considered for our model, it is not a concern as we use DL models, which are essentially black-box models and require explanations to understand the results. The explanations provided by our chosen methods of Grad-CAM, LIME, and SHAP also do not reveal any information about the architecture of the model itself; instead, they help us understand the regions of interest chosen by the model in each sample image. Each output of the models provides visual information that highlights the most important regions in the image.
@@ 5.2 Explanation Misuse
The factor of information misuse comes into consideration based on the information in the model that is leaked via explanations. A model can be stolen or replicated if considerable information about it can be gathered through the explanations. However, this factor does not apply to our model, as our methods of explanation do not leak any information regarding the architecture of the model.
@@ 5.3 Explanation Invarience
The third factor in the dimension of safety is explanation invariance. The primary objective of explanations is to develop an understanding of humans for each sample. Hence, an explainable system is expected to be stable and consistent.
The factor of stability defines that explanations of the same prediction by a model generated multiple times should result in the same explanation as well. This was evaluated in our model as well by implementing the three XAI techniques multiple times on the same image. An explanation generated by LIME results from the process of random perturbation, which generates simulated data in the surroundings of an instance for a simple linear classifier to predict and results in instability as it leads to the generation of different explanations for the same prediction [46]. Hence, while Grad-CAM and SHAP are stable and produce the same explanation for each prediction, the explanation provided by LIME is unstable, which has been shown in Figure 10.
Figure 10: Multiple LIME results for the same mammogram
The factor of consistency defines that explanations of a fixed model for similar data points should also be similar. We have evaluated the consistency of our model by using the ROI data provided in the original CBIS DDSM dataset and checking whether images with similar regions of interest have similar XAI results. The Figure 11 shows two similar ROI images with similar Grad-CAM results, which shows the consistency of Grad-CAM. SHAP works by assigning each feature an importance value for a particular prediction [47]. Hence, similar datapoints are assigned similar importance values, which shows the consistency of SHAP. The instability of LIME also makes LIME inconsistent as it generates different explanations for the same prediction; hence, comparison of similar data points does not produce similar explanations.
Figure 11: Consistency of Grad-CAM results (left column: ROI images, right column: Grad-CAM results)
@@ 5.4 Explanation Quality
The final safety measure is the explanation quality, which emphasizes the correctness of the explanation provided. While we have achieved the explainability of the model, it is necessary to analyze the correctness of the explanations provided by the different XAI functions to understand the performance of the black-box DL model. To analyze the explanations provided by XAI techniques for the performance of our model, the ROI images of the CBIS DDSM dataset have been used to compare the areas of importance highlighted by verified pathologists with the areas highlighted by XAI techniques. The ratio of similarity between the results of XAI and the original ROI images indicated how well the model is detecting the correct parts of the images for its classification. To create this ratio, we have created binary masks of the explanations provided by Grad-CAM and LIME which are shown in Figure 12.
@@
Figure 13 shows the original mammogram, the pre-processed image, the Grad-CAM output, the mask generated from the Grad-CAM heatmap, and the original ROI from the dataset.
@@
Figure 14 shows the original mammogram, the pre-processed image, the LIME output, the mask generated from the LIME segments, and the original ROI from the dataset.
The binary masks are then compared to the ROI images using Hausdorff distance. Hausdroff distance is a common measure for comparing the dissimilarity between image segmentation and point sets [48]. This method is commonly used in the medical domain to evaluate segmentation in medical images such as computed tomography (CT) and magnetic resonance imaging (MRI) [49]. The Hausdorff distance $\check{H}$ between two point sets $A$ and $B$ is the maximum distance between each point, as shown in the following equation [48].
@@
When comparing the ROI images with the masks produced by Grad-CAM, the Hausdorff distance values of our dataset range from 1 to 133, with the mean Hausdorff value being 18. When comparing the ROI images with the masks produced by Grad-CAM, the Hausdorff distance values of our dataset range from 1 to 160, with the mean Hausdorff value being 86. This indicates that the performance of the Grad-CAM explainability technique is better than the explainability provided by LIME.